a truck with a mass of 2200 kg travels at a constant velocity of 18 m/s, with a Fp=1400. What is the mu?
1
Expert's answer
2017-11-08T15:04:06-0500
For the constant velocity we need:
F_p=F_friction
F_friction=μmg.
μ=F_p/mg=1400/(2200∙9.8)=0.065.